OpenStudy (anonymous):

Help with question #1 below!

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

bubble 2

OpenStudy (anonymous):

How do you know?

OpenStudy (anonymous):

f= x^2+4 fof = (x^2+4)^2+4

OpenStudy (anonymous):

=x^4+8x^2+16+4

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

Thank you guys!